Please help me solve the following problem:
If one leg of a right triangle measures 8 feet and the the angle measures 30, what is the measure of the other leg?
The triangle looks like this:
30 degrees
--/!
-/-!x feet
/_!
8 feet
THANK YOU!!! =)

Answer by checkley79(3341)

What you have is a 30/60/90 triangle.
A^2+B^2=C^2
A=30 DEGREES, B=60 DEGREES.
This is a 1/2/sqrt3 triangle.
If A=8 Then: B=2*8=16 Ans. for the other leg.